OpenStudy (anonymous):

If the profits in your consulting business increase by 6% one year and decrease by 3% the following year, your profits are up by 3% over two years. ( true or false) If false - what % is your profit up? Help please.

OpenStudy (nurali):

That would be false. Your profits would be more because the 3% is coming off of a bigger sum.

OpenStudy (kropot72):

Let the reference profit be P. At the end of the first year the profit will be P * 1.06 At the end of the second year the profit will be (P * 1.06) * 0.97 = 1.0282P Therefore the percentage increase in profit over two years is \[\frac{1.0282P-P}{P} \times 100=you\ can\ calculate\]